for a special price one may choose three kind of deli meats, two kinds of cheese and one type of bread. if there are eight choices of meat, four choices of cheese, and five choices of bread, how many different sandwiches with three meats and two cheese will be possible?
Answer by stanbon(75887)

for a special price one may choose three kind of deli meats, two kinds of cheese and one type of bread. if there are eight choices of meat, four choices of cheese, and five choices of bread, how many different sandwiches with three meats and two cheese will be possible?
--------------------------------------------
# of ways to pick 3 meats: 8C3 = (8*7*6)/(1*2*3) = 56
# of ways to pick 2 cheese:4C2 = (4*3)/(1*2) = 6
# of ways to pick 1 bread: 5C1 = 5
-----
Total # of sandwiches: 56*6*5 = 1680
=====================================
Cheers,
Stan H.
=========